OpenStudy (anonymous):

At a pet store the total cost of 8 pounds of Brand X dog food and 1 pound of Brand Y dog food is $8.40 including tax. Total cost of 16 pounds of Brand X dog food and 8 pounds of Brand Y dog food is $24 including tax. What is the price per pound of Brand Y?

OpenStudy (radar):

Let x= price per lb of X Brand dog food Let y= price per lb of Y Brand dog food. Now set up your equations from the information you have provided. 8x + y = 8.40 16x + 8 y = 24.00 we will use the elimination process divide the 2nd equation by 8 and place the two equations so the can be subtracted. 8x + y = 8.40 2x + y = 3.00 (2nd equation divided by 8) subtract -------------------- 6x = 5.40 x=.90 or 90 cents substitute this value for x in the first equation above and get $1.20 for Brand y

OpenStudy (radar):

Any questions schulte?

OpenStudy (anonymous):

We substituted X for the first equation and got 10.2??

OpenStudy (radar):

Let me check 8x + y =8.40 substitute solution for x 8(.90) + y =8.40 substituting for x 7.20 + y = 8.40 solving for y y=8.40 - 7.20 how did you get 10.2?

OpenStudy (radar):

It looks like it is 1.20 a pound.

OpenStudy (anonymous):

Ya we got it Thanks!

OpenStudy (radar):

Good do you understand the elimination process?

OpenStudy (anonymous):

Yes we do!(:

OpenStudy (radar):

Good and good luck with your studies

OpenStudy (anonymous):

thank you ! you too